# T2M-HiFiGPT: Generating High Quality Human Motion from Textual Descriptions with Residual Discrete Representations

### 17 Dec 2023 v1

@@ 4.4 Comparison to state-of-the-art approaches
We show the comparison results in  table 3 and table 4 on HumanML3D [25] test set and KIT-ML [58] test set. From table 3, on both datasets, our reconstruction with RVQ-VAE reaches close performances to real motion and achieves better construction quality than the recent VQ-based methods, which suggests high-quality discrete representations learned by our RVQ-VAE. For the generation, our approach achieves state-of-the-art performance on the FID metric compared to the state-of-the-art method Att-T2M [81], while significantly surpasses them with the matching distance metric and recall precision. KIT-ML [58] and HumanML3D [25] are on different scales, which demonstrates the robustness of the proposed approach. Manually corrupting sequences and randomly dropping CLIP condition vectors during the training of double-tier GPT brings consistent improvement. A more detailed analysis is provided in the next section. Note that our framework can also automatically predict the generated motion length from the binary stop logits eq. 8, which is very intuitive and shown to be practical. Note that MDM [69], MotionDiffuse [80] and MLD [15] evaluate their models with the ground-truth motion length, which is not practical for real applications.
@@
We compare the computational time and network parameter quantity between our method and the most relevant GPT-based two-stage methods T2M-GPT [79] and Att-T2M [81] in table 1 and table 2. From table 1 and table 2, we can observe that at the reconstruction stage, all methods have almost the same parameter numbers. However, our RVQ-VAE has faster inference speed than the two VQ-VAE-based methods, while achieving the optimal reconstruction performance that is very close to the ground truth motion, confirming the effectiveness of our RVQ-VAE in the representation of human motion data. At the generation stage, our proposed double-tier GPT has four times fewer parameters and comparable inference speed while achieving SOTA performance in the text-to-motion task, which indicates our parameter efficacy and competitive computational cost in a practical system. Accordingly, our framework can both represent high-fidelity motion data and generate high-quality motion with textual description at a fast speed with fewer parameters.
@@
fig. 2 shows visual results on HumanML3D [25]. We compare our generations with the current state-of-the-art models: T2M-GPT [79], Att-T2M [81]. From the example in fig. 2, one can figure out that our model generates human motion with better text-to-motion consistency than the others. For example, the generated motions of Att-T2M [69] and T2M-GPT [79] either neglect some key actions such as ”stretches” and ”claps” or confuse the locomotion direction. Accordingly, they are not related to the semantics of the description. Note that more visualization results are provided in the supplementary material.

### 24 Dec 2023 v2

@@ 4.4 Comparison to state-of-the-art approaches
We show the comparison results in  table 3 and table 4 on HumanML3D [25] test set and KIT-ML [58] test set. From table 3, on both datasets, our reconstruction with RVQ-VAE reaches close performances to real motion and achieves better construction quality than the recent VQ-based methods, which suggests high-quality discrete representations learned by our RVQ-VAE. For the generation, our approach achieves state-of-the-art performance on the FID metric compared to the state-of-the-art GPT-based two-stage method Att-T2M [83] and diffusion-based method ReMoDiffuse [81] or FineMoGen [82], while significantly surpassing them with the matching distance metric and recall precision. KIT-ML [58] and HumanML3D [25] are on different scales, which demonstrates the robustness of the proposed approach. Manually corrupting sequences and randomly dropping CLIP condition vectors during the training of double-tier GPT brings consistent improvement. A more detailed analysis is provided in the next section. Note that our framework can both automatically predict the generated motion and its length from the binary stop logits eq. 8 in an online manner, which is very intuitive and demonstrated to be practical. In contrast, the diffusion-based methods such as MDM [69], FineMoGen [82], ReMoDiffuse [81], MotionDiffuse [80] and MLD [15] evaluate their models with the ground-truth motion length; their methods cannot automatically generate accurate motion length, which is very important for real applications.
@@
We compare the computational time and network parameter quantity between our method and the most relevant GPT-based two-stage methods, T2M-GPT [79] and Att-T2M [83] in table 1 and table 2. From table 1 and table 2, we can observe that at the reconstruction stage, all methods have almost the same parameter numbers. However, our RVQ-VAE has faster inference speed than the two VQ-VAE-based methods, while achieving optimal reconstruction performance that is very close to the ground truth motion, confirming the effectiveness of our RVQ-VAE in the representation of human motion data. At the generation stage, our proposed double-tier GPT has four times fewer parameters and comparable inference speed while achieving SOTA performance in the text-to-motion task, which indicates our parameter efficacy and competitive computational cost in a practical system. Accordingly, our framework can both represent high-fidelity motion data and generate high-quality motion with textual description at a fast speed with fewer parameters.
@@
fig. 2 shows some visual results on HumanML3D [25]. We compare our generations with the current GPT-based state-of-the-art models: T2M-GPT [79], Att-T2M [83] in fig. 3. From the examples in fig. 3, one can figure out that our model generates human motion with better text-to-motion consistency than the others. For example, the generated motions of Att-T2M [69] and T2M-GPT [79] either neglect some key actions such as ”stretches” and ”claps” or confuse the locomotion direction. Accordingly, they are not related to the semantics of the description. Note that more visualization results are provided in the supplementary material.
Figure 2: Visual results on HumanML3D [25]. Our T2M-HiFiGPT can generate precise and high-fidelity human motion consistent with the given text descriptions.